"Find all real solutions to each equation or inequality. For the
inequalities, also sketch the graph of the solution set.
1/x - 1/(x-1)= -1/6 Thank you."

Answer by nerdybill(7384)

1/x - 1/(x-1)= -1/6
.
Start by multiplying both sides with a common denominator: 6(x)(x-1)
.
6(x)(x-1)[1/x - 1/(x-1)]= 6(x)(x-1)[-1/6]
6(x-1) - 6(x) = -(x)(x-1)
6x - 6 - 6x = -x^2 + 1
- 6 = -x^2 + 1
- 6 + x^2 = 1
x^2 = 7
x = (+-)sqrt(7)